OpenStudy (anonymous):

Factor by finding the difference of two perfect squares: (121m^6-p^16)

OpenStudy (raden):

121m^6 = (11m^3)^2 and p^16 = (p^8)^2 the difference of two perfect squares like : a^2 - b^2 = (a+b)(a-b) so, 121m^6 - p^16 can be facored as = (11m^3)^2 - (p^8)^2 = (11m^3 + p^8)(11m^3 - p^8)